During the year ended December 31, 2018, Kellys Camera Shop had sales revenue of $190,000, of which $95,000 was on credit. AGive the required journal entries for the two events in December. (If no entry is required for a transaction/event, select NGive the required journal entries for the two events in December. (If no entry is required for a transaction/event, select N
Answer #1
1
Dec 10,2018 Allowance for Doubtful Accounts 1700
         Accounts Receivable 1700
Dec 31,2018 Bad Debts expense 1900 =95000*2%
      Allowance for Doubtful Accounts 1900
2A
Balance Sheet(Partial)
At December 31,2018
Current Assets
Accounts Receivable 32300 =11000+95000-72000-1700
Allowance for Doubtful Accounts 840 =640-1700+1900
Accounts Receivable, net of allowance 31460
2B
Income Statement(Partial)
Year ending December 31,2018
Operating expenses:
Bad Debts expense 1900
3
Yes, the rate seems to be reasonable.